OpenStudy (anonymous):

HELLP!!! What's the derivative (F'(x) ) of F(x) (√x-3+x^2)/4x ??

OpenStudy (anonymous):

\[f(x)=\frac{ \sqrt{x^2+x-3} }{ 4x }\] is this right? your equation is a little vague...

OpenStudy (anonymous):

no the radical is just for the second x

OpenStudy (anonymous):

\[\frac{\sqrt{x}-3+x^2}{4x}\]?

OpenStudy (anonymous):

or is the \(-3\) under the radical as well? it makes a huge difference in the method

OpenStudy (anonymous):

I think the -3 is under the radical as well

OpenStudy (anonymous):

guess i'll never know

OpenStudy (anonymous):

Yes like you wrote it! :)

OpenStudy (anonymous):

split it up into \[\frac{ \sqrt{x} }{ 4x } - \frac{3}{4x} + .25x\]

OpenStudy (anonymous):

ok thanks :)